Two angles are supplementary. One angle measures 15 degrees more than 4 times the other. what is the measure of the larger angle in degrees?
Answer by macston(5194)

.
S=smaller angle; L=larger angle=4S+15 degrees
.
S+L=180 degrees Substitute for L.
S+4S+15 degrees=180 degrees Subtract 15 degrees from each side.
5S=165 degrees Divide each side by 5.
S=33 degrees The smaller angle is 33 degrees.
.
L=4S+15 degrees=4(33 degrees)+15 degrees=132 degrees+15 degrees=147 degrees
ANSWER: The larger angle is 147 degrees.